OpenStudy (nuccioreggie):

For a class pizza party, Esther bought ten bottles of juice and 9 bottles of water. Let j represent the cost of juice and w represent the cost of water. Write an expression to determine the total cost for the drinks. Then, calculate the cost if juice costs two dollars and water costs one dollar. ten j plus nine w; twenty-nine dollars ten j plus nine w; twenty-eight dollars nine j plus ten w; twenty-eight dollars nine j plus ten w; twenty-nine dollars

OpenStudy (anonymous):

10j x 2 = 20.00 1 x 9w= 9.00 20.00 +9.00 =?

OpenStudy (anonymous):

what would it be to you?

OpenStudy (anonymous):

its pretty obvious

OpenStudy (triciaal):

10 j + 9 w = cost of drinks

OpenStudy (anonymous):

@nuccioreggie

OpenStudy (triciaal):

find when j = 2 and w = 1

OpenStudy (nuccioreggie):

b ?

OpenStudy (triciaal):

10 j + 9 w = cost j = 2 and w = 1 10 * 2 + 9* 1 = 20 + 9 = 29 answer choice a